OpenStudy (anonymous):

where does the function intersect the x-axis ? y=x^2-2x-15

OpenStudy (anonymous):

it will intersect the x axis when y = 0. Substitute y=0 into the equation and then solve.

OpenStudy (anonymous):

what two numbers add up to 2 and multiply to -15 ??

OpenStudy (anonymous):

-3 and 5

OpenStudy (anonymous):

so.. 0=(x-3)(x+5) ^x=3 ^x=-5 then 0=0

OpenStudy (anonymous):

setting 0 for x, right ?

OpenStudy (anonymous):

yup. So the graph intersects the x-axis at x=3 and -5

OpenStudy (anonymous):

oh thank you so much. i understand now(:

OpenStudy (anonymous):

No problem. Hopefully you won't stink at math anymore. lol

OpenStudy (anonymous):

haha! thats a good one. lol